OpenStudy (anonymous):

how do you expand and simplfy these kind of questions, 3c (c - d)

OpenStudy (anonymous):

do you know how to distribute? a(c+d) = ac + ad

OpenStudy (anonymous):

no all i know is do to the more simple one where it is like 3(d + 5)

OpenStudy (anonymous):

what does that one you wrote equal to?

OpenStudy (anonymous):

whish is 3 x d=d3 3x5=15 = d3+15

OpenStudy (anonymous):

good now what about this simple one a(b+2)?

OpenStudy (anonymous):

a x b= ab i think :/ a x 2= a2 ab=a2

OpenStudy (anonymous):

its not equals though its a(b+2) = ab + 2a, which is what you got right?

OpenStudy (anonymous):

i got ab and a2 so yh sorry i was gettin somethin

OpenStudy (anonymous):

ok now a(c-d)?